Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#265 closed flight processing (fixed)

GB08/08, flight day 252b/2009, Wytham woods

Reported by: mark1 Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_252b_Wytham_woods

Data arrived from ARSF via SATA disk L on 7th October 2009.

Scientific objective: see ~/arsf/arsf_data/in_progress/2008/ARSF_Applications/GB_and_NL/GB08-08 Science Case.pdf

Priority: alpha-4-High

PI: Richard Harding

Sensors:

  • Eagle (30/04/2010)
  • Hawk (30/04/2010)
  • Leica LIDAR (25/05/2010)
  • RCD images (12/02/2010)

Change History (38)

comment:1 Changed 10 years ago by mark1

Still awaiting flight logsheet for this project.

Also, DEM creation gave errors when reading in tiles to grass but still output a DEM. Check it is OK when processing.

comment:2 Changed 10 years ago by mark1

Have moved original eagle files into boresight_252a and moved the eagle files from boresight_252a into this project.

comment:3 Changed 10 years ago by mark1

  • Milestone set to 2009 Data processing completion

comment:4 Changed 10 years ago by mark1

logsheet arrived and unpacked

comment:5 Changed 9 years ago by mggr

Re: #281 (Hawk frame rates are double what they should be), the following lines are affected and the .hdr file will need to be edited to halve the fps number:

arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-10.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-11.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-12.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-13.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-2.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-3.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-4.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-6.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-7.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-8.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIR252a_09-9.hdr: predicted length half recorded length, frame rate doubling problem?
arsf_internal/boresight_2009-252a_little_riss/hawk/SWIRdataCube.hdr: predicted length half recorded length, frame rate doubling problem?

comment:6 Changed 9 years ago by mark1

Photograph delivery created and checked. No log file to tag images with.

comment:7 Changed 9 years ago by mark1

  • Description modified (diff)

Photography data dispatched to PI on 12/02/2010

comment:8 Changed 9 years ago by emca

  • Owner set to emca
  • Status changed from new to assigned

comment:9 Changed 9 years ago by knpa

  • Other processors set to knpa

Started Hyperspectral Processing

comment:10 Changed 9 years ago by knpa

Applanix already processed by emca and .cfg file generated.

emca recommends regenerating .cfg file as the script has been updated (and needs boresight parameters - see below).

comment:11 Changed 9 years ago by knpa

This flight requires someone to do the calibration of day 252 before it can be processed.

comment:12 Changed 9 years ago by knpa

Calibration complete, restarting project.

comment:13 Changed 9 years ago by knpa

correct dem is sp40sp50.dem

comment:14 Changed 9 years ago by knpa

Hawk suffering from dropped bands issue, have halved frame rate from 30 to 15 fps.

comment:15 Changed 9 years ago by knpa

Dem did not enclose line 2, have extended to sp40sp41sp50sp51.

comment:16 Changed 9 years ago by knpa

Sync error on eagle line 1.

comment:17 Changed 9 years ago by knpa

Hawk line 2 has only 17M of data, and is in wrong place. Don't think it is salvageable.

comment:18 Changed 9 years ago by knpa

Correction - above is referring to Hawk line 1. Hawk line 2 is fine.

comment:19 Changed 9 years ago by knpa

SCTs

e h
1 1.18 (sync) no data
2 0.02 -0.03
3 0.03 -0.02
4 0.00 -0.07
5 0.00 -0.07
6 0.02 0.02
7 0.01 -0.03

comment:20 Changed 9 years ago by knpa

PI on logsheet is given as Andrew Wilson, but ticket and application form and emca all confirm it is Richard Harding.

comment:21 Changed 9 years ago by knpa

Hyperspectral ready to be delivery checked (29/0410)

comment:22 Changed 9 years ago by mark1

Delivery checked: Example commands tested and work OK. Minor corrections made to readme. All data viewed and looks good.

Before delivery should change PI on log sheet to the correct one. Otherwise ready to deliver.

comment:23 Changed 9 years ago by knpa

PI changed on logsheet, old saved as logsheet_original.xxx

comment:24 Changed 9 years ago by knpa

HYPERSPECTRAL DELIVERED 30/04/10

comment:25 Changed 9 years ago by knpa

  • Description modified (diff)

comment:26 Changed 9 years ago by emca

LIDAR proccessing complete.
Ready for delivery check at
~airborne/workspace/GB08_08-2009_252b_Wytham_woods/delivery/20100514/GB08_08

comment:27 Changed 9 years ago by knpa

  • Other processors knpa deleted

comment:28 Changed 9 years ago by mark1

LIDAR delivery checked and everything looks good.
Ready to deliver.

comment:29 Changed 9 years ago by emca

LIDAR data sent to PI 25/05/2010

comment:30 Changed 9 years ago by emca

  • Description modified (diff)

comment:31 Changed 9 years ago by emca

LIDAR and Hyperspectral data has been rsync'd to
/users/rsg/arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_252b_Wytham_woods

Currently not enough space for the photography delivery

comment:32 Changed 9 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from emca to knpa

Project appears to have been completely r-synced now, beginning archiving.

comment:33 Changed 9 years ago by knpa

No rcd logs present for this project, therefore no event file in delivery either.

comment:34 Changed 9 years ago by knpa

No screenshot with vectors in lidar delivery.

comment:35 Changed 9 years ago by knpa

Made available as tarball to NEODC.

comment:36 Changed 9 years ago by knpa

Download and backed-up by NEODC.

Moved to archive disk.

comment:37 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleting workspace version and closing ticket.

comment:38 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.